OpenStudy (anonymous):

Convert 86000 cm to km

OpenStudy (anonymous):

.86 km

OpenStudy (anonymous):

100 cm = 1m 1000 m = 1000 (100)cm = 1 km 86000 cm = ? km

OpenStudy (alexandervonhumboldt2):

i'm form russia where we use metric i will hlp so 86000/(1000*100)

OpenStudy (anonymous):

Thank you, 100 cm = 1m 1000 m = 1000 (100)cm = 1 km 86000 cm = .86

OpenStudy (alexandervonhumboldt2):

0.86

OpenStudy (michele_laino):

since 1Km=10^5 cm, you have to divide by 10^5, so: \[\frac{ 86000 }{ 10^{5} }=0.86\]

OpenStudy (alexandervonhumboldt2):

@Starbucks show your steps

OpenStudy (alexandervonhumboldt2):

dont gve only answer

OpenStudy (anonymous):

Google is my steps . Okay ? okay . (:

OpenStudy (anonymous):

@Starbucks thanks for the useful help :)

OpenStudy (anonymous):

Your welcome . (: